# Directional and contra-directional coupling in Huygens’ metawaveguide microring resonators

### v1

@@ 3 Conclusion
Funding The project was funded by the (1) National Research Council Canada (NRC)’s (i) Postdoctoral Fellowship Program, (ii) Collaborative Science, Technology and Innovation Program (CSTIP) Small Teams (ST-R2-01-02), and by the (2) Deutsche Forschungsgemeinschaft (DFG, German Research Foundation, project number 448835038).
Acknowledgment M. Saad Bin-Alam acknowledges the support from the NRC’s Postdoctoral Fellowship in Metamaterial Integrated Photonics, and is thankful to Dr. Shurui Wang and Martin Vachon for their help in the experimental characterizations.
Disclosures The authors declare no conflicts of interest.
Data Availability Statement Data underlying the results presented in this paper are not publicly available at this time but may be obtained from the authors upon reasonable request.
Supplemental document See Supplement 1 for supporting content.

### v2

@@ Conclusion
Acknowledgements
The project was funded by the (1) National Research Council Canada (NRC)’s (i) Postdoctoral Fellowship Program, (ii) Collaborative Science, Technology and Innovation Program (CSTIP) Small Teams (ST-R2-01-02), and by the (2) Deutsche Forschungsgemeinschaft (DFG, German Research Foundation, project number 448835038).
M. Saad Bin-Alam acknowledges the support from the NRC’s Postdoctoral Fellowship in Metamaterial Integrated Photonics, and thanks Dr. Shurui Wang and Martin Vachon for their help in the experimental characterizations.
Supporting information
The Supporting Information is available free of charge.
Multipole expansion and decomposition analysis; Forward directional cross-coupling; Resonance spectra, free spectral range (FSR), group index and group velocity dispersion; Resonance splitting and linewidth narrowing; Characteristics of the contra-directional coupler (CDC)-loaded racetrack resonators